OpenStudy (anonymous):

a recipe that makes 6 serving calls for 1/3 cups of suger.arlean modifes the recipe so that it can serve 9. How many cups of suger does arlean need

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

you take 6 servings and divide by 1/3 and you will find how much sugar per cup then you multiply by 9

OpenStudy (anonymous):

0.45? am i correct

OpenStudy (anonymous):

\[1/3 \approx 0.3repeating\]

OpenStudy (anonymous):

so the answer for the question will be 1/3

OpenStudy (anonymous):

@hartnn can help

OpenStudy (anonymous):

(rasing) my hand sombody)

OpenStudy (anonymous):

sombody here

OpenStudy (anonymous):

i got ya. i think of it like this. 6 servings equals 1/3 of a cup. 3 servings more is an extra half serving of sugar. OR 1/3/2 = 1/6 1/3 + 1/6. find the gcf and add them up

OpenStudy (tkhunny):

6 servings : 1/3 cup Find the unit amount per serving. This is what Zie is trying to tell you Divide by 6 6/6 = 1 serving : (1/3)/6 = 1/18 cup Find the required amount. Multiply by 9 9*1 = 9 servings : 9*(1/18) = 1/2 cup Alternatively, use the phrase "half again as much". This is 150% or 1.5 Multiply by 1.5 6*1.5 = 9 servings : (1/3)*1.5 = (1/3)(3/2) = 1/2 cup

OpenStudy (anonymous):

can you help that kind of makes no scence on how to do it

OpenStudy (anonymous):

ohhh i get it can we do another one do we have time

OpenStudy (anonymous):

A recipe for 1 lb of trail mix calls for 3.2 oz of cashews. How many ounces of cashews are needed to make 4 lb of trail mix?

OpenStudy (anonymous):

same concept. What you do to one side, you do to the other. 1lb =3.2oz you want 4lbs so.. 1*4 lbs = 3.2 *4 oz cashews

OpenStudy (anonymous):

do i do 3.2*1=3.2/4=0.8 oh i did it wrong so then once i do 3.2*4 that will be the answer

OpenStudy (anonymous):

right

OpenStudy (tkhunny):

In the second case, you already have the unit amount. There is no need to find it!!

OpenStudy (anonymous):

ok so then 12.8oz is the answer

OpenStudy (anonymous):

right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Last month, Edwin had $2325.35 that he divided into two bank accounts. He deposited into a savings account. He deposited the rest into a checking account. How much money did Edwin deposit into his checking account?

OpenStudy (anonymous):

he deposited 2/5

OpenStudy (tkhunny):

Okay, time to raise my hand and say, "Hold on there!" Show your work. You have had examples enough. Generally, post new problem on new threads.

OpenStudy (anonymous):

i need help fast before i leave thats why and sorry

OpenStudy (tkhunny):

No excuses. Show your work.

OpenStudy (anonymous):

i am not lying i have to go to the dentisest i will give you my phone numder so you can talk to my mom if you dont belive me

OpenStudy (anonymous):

just believe me nobody in my family beliave me for anything

OpenStudy (anonymous):

Last month, Edwin had $2325.35 that he divided into two bank accounts. He deposited 2/5 into a savings account. He deposited the rest into a checking account. How much money did Edwin deposit into his checking account?

OpenStudy (anonymous):

checking account is going to be the whole amount minus what was deposited into the savings. so...

OpenStudy (anonymous):